OpenStudy (anonymous):

Which of the following is in the solution set for the inequality –6 > n – 4?

OpenStudy (anonymous):

Add 4 to both sides in order to get 'n' by itself. Now you'll end up with -2>n so n is less than -2.

OpenStudy (anonymous):

thanks @jam333 can you help with another prolm?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

solve the inequality s-12>_-1

OpenStudy (anonymous):

@jam333

OpenStudy (anonymous):

Is that s-12>-1 or s-12>=-1. If you are typing 'greater than or equal to' it's this >=

OpenStudy (anonymous):

no its that symbol with the line under it

OpenStudy (anonymous):

Yes, that symbol with a line under it means 'greater than or equal to' and we type it as >= because we can't type the other symbol with a regular keyboard. s-12>=-1 You want the 's' alone so you add 12 to both sides of the inequality to get s >= 11. That's the answer. S is greater than or equal to 11.

OpenStudy (anonymous):

It's kind of confusing to add one number to a negative number. Think of it as temperature. Let's say someone said it was -1 degrees outside. Then they said it got 4 degrees hotter. You'd have to add 4 to -1. So -1 + 4 is 3. So now it's 2 degrees. I like to picture looking at a thermometer going up and down. If it's -1 degrees and it go 5 degrees colder, then you'd do -1 - 5 = -6.